Current of 50A flows to east on stationery wire . A second wire is 1cm below it . How much current must be flowing in the second wire so that doesn't fall due to gravity ? What is the direction of the current ? The mass per unit length of the second wire is 5g/m

1
Expert's answer
2022-04-26T15:19:30-0400

"mg=\frac{\mu_0}{2\pi}\frac{I_1I_2l}d,"

"I_2=\frac{2\pi mgd}{\mu_0 I_1 l}=50~A."